OpenStudy (anonymous):

What is the sum of the arithmetic sequence 174, 168, 162 …, if there are 37 terms

OpenStudy (anonymous):

See, a = 174 d = 168 - 174 = -6 Find the last term Using: \[\large t_n = a + (n-1)d\] Here n= 37.. So, \[\large t_n = 174 + (37-1)(-6)\] Solve this for \(t_n\),,

OpenStudy (anonymous):

-42?